# The GD-1 stellar stream perturber as a core-collapsed self-interacting dark matter halo

### v1

@@ 1 Introduction
In this work, we assume that the GD-1 perturber is a dark matter subhalo and explore its formation in within the framework of self-interacting dark matter (SIDM), see Tulin & Yu (2018); Adhikari et al. (2022) for reviews and references therein. The gravothermal evolution of an SIDM halo occurs in two sequential phases. In the core-forming phase, dark matter self-interactions transport heat inward, resulting in a shallow density core, while in the core-collapsing phase, heat transfer reverses, leading to a higher central density than in the CDM counterpart (e.g. Balberg et al., 2002; Koda & Shapiro, 2011; Essig et al., 2019; Feng et al., 2021). Notably, SIDM models with large cross sections could explain the high density of the strong lensing perturber for SDSSJ0946+1006 (Nadler et al., 2023) and the low density of the Crater II satellite galaxy (Zhang et al., 2024), both challenging CDM. It is intriguing to explore the SIDM scenario to account for the high density of the GD-1 perturber.
@@
The rest of this paper is organized as follows: In Sec. 2, we discuss the properties of CDM halos in the cosmological zoom-in simulation of a Milky Way analog. In Sec 3, we introduce the setup of our N-body simulations. In Sec. 4, we present the properties of our simulated SIDM and CDM subhalos and compare them with the GD-1 perturber. In Sec. 5, we discuss future investigations for further improvement and conclude. In Appendix A, we present the SIDM simulation of an isolated halo for testing numerical artifacts that could lead to violation of energy conservation. In Appendix B, we show the convergence test.
@@ 2 CDM halos of a Milky Way analog
We first present progenitor halos from a cosmological zoom-in CDM-only simulation of a Milky Way analog (Yang et al., 2023a; Nadler et al., 2020b), with initial conditions drawn from the suite in Mao et al. (2015). This simulated system includes a main halo with a mass of $1.6\times 10^{12}~{}M_{\odot}$ and a Large Magellanic Cloud analog. The simulation has a particle mass of $4\times 10^{4}~{}M_{\odot}$, a Plummer-equivalent softening length of $\epsilon=0.08~{}{\rm kpc}~{}h^{-1}$, and a spline length of $\ell=2.8\epsilon=0.22~{}{\rm kpc}~{}h^{-1}$.
@@
We select subhalos of the main halo with the virial mass larger than $10^{8}~{}M_{\odot}$ at $z=0$, and then identify their progenitors at infall. With the mass cut, there will be at least $2500$ simulation particles for each progenitor halo so that we can accurately reconstruct its density profile. The radial resolution of the cosmological simulation $\ell\approx 0.3~{}{\rm kpc}$ is more than an order of magnitude larger than the radial scale relevant for the GD-1 perturber. To overcome this resolution limit, we fit each progenitor halo with a Navarro-Frenk-White (NFW) profile (Navarro et al., 1997) for the region $r>0.3~{}{\rm kpc}$,
$\rho(r)=\frac{\rho_{s}}{\left(r/r_{s}\right)\left(1+r/r_{s}\right)^{2}},$ (1)
where $r_{s}$ and $\rho_{s}$ are the scale radius and density, respectively. We determine the $r_{s}$ and $\rho_{s}$ values for each progenitor at infall and the overall fit quality is excellent. We then extrapolate the NFW profile inward and compute the total enclosed mass within $r=10~{}{\rm pc}$.
@@
Figure 1 (left) shows the density profiles for the $125$ progenitor halos at infall (blue). We also present an NFW fit to one of the progenitors (black), which will be used as the initial condition for our SIDM simulations; see detailed comparison in the inset panel. The simulated density profile is flattened for $r\lesssim 0.3~{}{\rm kpc}$ due to the resolution limit. However, we expect that the NFW profile provides a good approximation for extrapolating the density inward before the halo undergoes significant tidal stripping.
@@
Figure 1 (middle) shows the enclosed mass within $10~{}{\rm pc}$ vs virial mass of the progenitor halos at infall. For comparison, we include a reference case from the viable parameter region of the GD-1 perturber in Bonaca et al. 2019 (their Figure 6): a Hernquist scale radius of $r_{\rm H}=15~{}{\rm pc}$ and a total mass of $M=4.6\times 10^{5}~{}M_{\odot}$, which approximately corresponds to a substructure with the minimum density required to explain the spur and gap features of the GD-1 stream. For this reference case, the enclosed mass within $10~{}{\rm pc}$ is $\approx 7.4\times 10^{4}~{}M_{\odot}$, as denoted by the horizontal line in the middle panel. We see that none of the CDM progenitor halos are sufficiently dense to be the perturber, and this conclusion holds when comparing the enclosed mass within $r=15~{}{\rm kpc}$. It is important to note that the inner density of these CDM halos would further decrease as they evolve within the Milky Way’s tidal field.
@@
where $C=0.75$ is a numerical factor and $\sigma_{\mathrm{eff}}$ is the effective cross section (Yang & Yu, 2022). For simplicity, we assume a constant cross section in this work. Figure 1 (right) shows the collapse time vs virial mass for the progenitors, where we have taken $\sigma/m=50~{}{\rm cm^{2}~{}g^{-1}}$. There are five halos that could collapse within $10~{}{\rm Gyr}$. Since tidal stripping could speed up the onset of the collapse (Nishikawa et al., 2020; Sameie et al., 2020; Kahlhoefer et al., 2019; Yang & Yu, 2021; Zeng et al., 2022), we expect that more halos would be in the collapse phase after they evolve in the tidal field and their overall mass would be reduced as well.
@@ 3.2 The Milky Way model
These parameters are motivated by the Milky Way mass model in McMillan (2016). Note that the stellar bulge and disk profiles are slightly different from those in McMillan (2016), but we have checked that the difference in the potential is within a few percent in the regions $\sqrt{R^{2}+z^{2}}>10~{}{\rm kpc}$, where are most relevant for our simulated subhalo. Since the host halo is treated as a static potential, we neglect dark matter particle scatterings between the host halo and the subhalo. This approximation is well justified for velocity-dependent SIDM models with $\sigma/m\lesssim 1~{}{\rm cm^{2}~{}g^{-1}}$ at $V_{\rm max}\sim 200~{}{\rm km~{}s^{-1}}$ (Nadler et al., 2020a).
@@ 3.3 Orbital parameters
Figure 2: Left: evolution of the enclosed mass within $10~{}{\rm pc}$ for the CDM (blue), SIDM30 (amber), SIDM50 (orange), and SIDM100 (pink) subhalos. The horizontal dashed gray line denotes the inner mass within $10~{}{\rm pc}$ of the reference Hernquist profile, as shown in Figure 1 (middle). Middle: evolution of the bound mass for the simulated CDM and SIDM subhalos. Right: corresponding density profiles for the simulated CDM and SIDM subhalos at $t=10~{}{\rm Gyr}$, along with the initial NFW profile (black). The dotted blue line denotes a reconstructed density profile for the CDM subhalo using an analytical function proposed by Errani & Navarro (2021). The gray shaded region denotes the viable range for the GD-1 perturber, converted from Figure 6 of Bonaca et al. (2019), while the dashed gray line represents the reference Hernquist profile.
@@ 4 Results
Figure 2 (left) shows the evolution of the enclosed mass within the inner $r=10~{}{\rm pc}$ for CDM (blue), SIDM30 (amber), SIDM50 (orange), and SIDM100 (pink) subhalos. For CDM, the inner mass decreases monotonically due to tidal stripping. In contrast, for SIDM, the mass initially decreases sharply due to core expansion, followed by an increase as core collapse occurs. By $t\approx 10~{}{\rm Gyr}$, the inner mass of the SIDM subhalos is an order of magnitude higher than that of the CDM subhalo, aligning well with the reference Hernquist profile (horizontal line). Additionally, the collapse times are $t_{c}\sim 6~{}{\rm Gyr},~{}4~{}{\rm Gyr}$, and $2~{}{\rm Gyr}$ for SIDM30, SIDM50, and SIDM100, respectively, about a factor of two shorter than those estimated using Equation 2, which is calibrated for isolated halos. For subhalos, tidal stripping can accelerate the onset of gravothermal collapse.
@@
We note that the CDM subhalo has a small density core near the center. This is due to the resolution limit as $\ell=2.8\epsilon\approx 5.6~{}{\rm pc}$, although the simulated subhalo contains more than $6\times 10^{5}$ simulation particles at $t=10~{}{\rm Gyr}$. We fit the density profile using the analytical function $\rho(r)=\rho_{\rm cut}\exp(-r/r_{\rm cut})r_{\rm cut}/r$ from Errani & Navarro (2021), which is proposed to model a tidally-stripped CDM halo. With $\rho_{\rm cut}\approx 1.2\times 10^{8}~{}M_{\odot}~{}{\rm kpc^{-3}}$ and $r_{\rm cut}\approx 0.22~{}{\rm kpc}$, we find a good fit for the region $r\gtrsim 10~{}{\rm pc}$, see Figure 2 (right, dotted blue). The fitted function has a cusp $\rho(r)\propto r^{-1}$ near the center and it provides a correction to the core due to the resolution limit. For the fitted profile, the enclosed mass within $10~{}{\rm pc}$ is $1.6\times 10^{4}~{}M_{\odot}$. However, even with this correction, the CDM subhalo remains insufficiently dense to explain the high density of the GD-1 perturber.
@@ 5 Discussions and conclusion
The subhalo we used to demonstrate the SIDM scenario for the GD-1 perturber has an infall mass of $\approx 3\times 10^{8}~{}M_{\odot}$. Interestingly, this is near the upper limit on the peak mass of subhalos that host currently-observed satellite galaxies in the Milky Way  (Jethwa et al., 2018; Nadler et al., 2020b). Thus, it remains an open question whether the perturber is a truly dark substructure, devoid of a galaxy. It would be interesting to conduct similar simulations with live stellar particles and investigate the detectability of such substructures. This is timely as the Rubin Observatory (Ivezić et al., 2019; Drlica-Wagner et al., 2019) is expected to discover more faint satellite galaxies in the near future. Furthermore, our scenario should also apply to smaller infall masses below $\sim 10^{8}~{}M_{\odot}$. Indeed, for SIDM models with large velocity-dependent cross sections, the population of core-collapsing (sub)halos increases as the mass decreases (e.g. Yang et al., 2023a; Nadler et al., 2023). Thus stellar streams like GD-1 can probe both population and density profile of core-collapsing subhalos even below the mass threshold for galaxy formation.

### v2

@@ 1 Introduction
In this work, we assume that the GD-1 perturber is a dark matter subhalo and explore its formation in within the framework of self-interacting dark matter (SIDM), see Tulin & Yu (2018); Adhikari et al. (2022) for reviews and references therein. The gravothermal evolution of an SIDM halo occurs in two sequential phases. In the core-forming phase, dark matter self-interactions transport heat inward, resulting in a shallow density core, while in the core-collapsing phase, heat transfer reverses, leading to a higher central density than in the CDM counterpart (e.g. Balberg et al., 2002; Koda & Shapiro, 2011; Essig et al., 2019; Feng et al., 2021). Notably, SIDM models with large cross sections could explain the high density of the strong lensing perturber for SDSSJ0946+1006 (Nadler et al., 2023; Minor et al., 2021; Vegetti et al., 2010) and the low density of the Crater II satellite galaxy (Zhang et al., 2024; Borukhovetskaya et al., 2022), both challenging CDM. It is intriguing to explore the SIDM scenario to account for the high density of the GD-1 perturber.
@@
The rest of this Letter is organized as follows: In Section 2, we discuss the properties of CDM halos in the cosmological zoom-in simulation of a Milky Way analog. In Section 3, we introduce the setup of our N-body simulations. In Section 4, we present the properties of our simulated SIDM and CDM subhalos and compare them with the GD-1 perturber. In Section 5, we discuss future investigations for further improvement and conclude. In Appendix A, we present the SIDM simulation of an isolated halo for testing numerical artifacts that could lead to violation of energy conservation. In Appendix B, we show the convergence test.
@@ 2 CDM halos of a Milky Way analog
We first present progenitor halos from a cosmological zoom-in CDM-only simulation of a Milky Way analog (Yang et al., 2023a; Nadler et al., 2020b), with initial conditions drawn from the suite in Mao et al. (2015). This simulated system includes a main halo with a mass of $1.14\times 10^{12}~{}M_{\odot}~{}h^{-1}\approx 1.6\times 10^{12}~{}M_{\odot}~{}(h=0.7)$, and a Large Magellanic Cloud analog. The simulation has a particle mass of $4\times 10^{4}~{}M_{\odot}~{}h^{-1}$, a Plummer-equivalent softening length of $\epsilon=0.08~{}{\rm kpc}~{}h^{-1}$, and a spline length of $\ell=2.8\epsilon=0.22~{}{\rm kpc}~{}h^{-1}$, the characteristic length scale of the smoothing kernel used to calculate gravitational forces between particles (Springel et al., 2001; Springel, 2005).
@@
We select subhalos of the main halo with the virial mass larger than $10^{8}~{}M_{\odot}~{}h^{-1}$ at $z=0$, and then identify their progenitors at infall. With the mass cut, there will be at least $2500$ simulation particles for each progenitor halo so that we can accurately reconstruct its density profile. The radial resolution of the cosmological simulation $\ell\approx 0.3~{}{\rm kpc}$ is more than an order of magnitude larger than the radial scale relevant for the GD-1 perturber. To overcome this resolution limit, we fit each progenitor halo with a truncated Navarro-Frenk-White (NFW) profile (Errani & Navarro, 2021) for the region $r>0.3~{}{\rm kpc}$,
$\rho(r)=\frac{\rho_{s}}{\left(r/r_{s}\right)\left(1+r/r_{s}\right)^{2}}\times\frac{\exp(-r/r_{\rm cut})}{(1+r_{s}/r_{\rm cut})^{0.3}},$ (1)
where $\rho_{s}$ and $r_{s}$ are the scale density and radius, respectively, and $r_{\rm cut}$ is the truncation radius due to tidal stripping. We determine the three parameters for each progenitor at infall, achieving excellent overall fit quality. The truncated NFW profile is then extrapolated inward to compute the total enclosed mass within $r=10~{}{\rm pc}$. Additionally, we confirm that many progenitor halos can be well-fitted with the standard NFW profile, while some exhibit density profiles slightly steeper than $r^{-3}$ in the outer regions. For these cases, the standard NFW fit may introduce bias and overestimate the central density. However, the truncated NFW profile provides a significantly better fit.
@@
Figure 1 (left) shows the density profiles for the $125$ progenitor halos at infall (blue). We also present the fit to one of the progenitors (black), which will be used as the initial condition for our SIDM simulations; see the detailed comparison in the inset panel. For this halo, the standard NFW profile provides a good fit. The simulated density profile is flattened for $r\lesssim 0.3~{}{\rm kpc}$ due to the resolution limit. However, we expect that the NFW profile provides a good approximation for extrapolating the density inward before the halo undergoes significant tidal stripping.
@@
Figure 1 (middle) shows the enclosed mass within $10~{}{\rm pc}$ vs virial mass of the progenitor halos at infall. For comparison, we include a reference case from the viable parameter region of the GD-1 perturber in Bonaca et al. 2019 (their Figure 6): a Hernquist scale radius of $r_{\rm H}=15~{}{\rm pc}$ and a total mass of $M=4.6\times 10^{5}~{}M_{\odot}$, which approximately corresponds to a substructure with the minimum density required to explain the spur and gap features of the GD-1 stream. For this reference case, the enclosed mass within $10~{}{\rm pc}$ is $\approx 7.4\times 10^{4}~{}M_{\odot}$, as denoted by the horizontal line in the middle panel. We see that none of the CDM progenitor halos are sufficiently dense to be the perturber, and this conclusion holds when comparing the enclosed mass within $r=15~{}{\rm pc}$. The inner density of these CDM halos would further decrease as they evolve within the Milky Way’s tidal field. The progenitor CDM halos shown in Figure 1 correspond to subhalos with masses $>10^{8}~{}M_{\odot}~{}h^{-1}$ at $z=0$. We plan to relax this mass threshold and examine halos with lower masses. Based on the resolution limit in the cosmological CDM simulation (Yang et al., 2023a), we expect to reconstruct the density profiles of progenitors for subhalos with masses a few times $10^{7}~{}M_{\odot}~{}h^{-1}$ using the truncated NFW profile in Equation 1. A more detailed investigation will be deferred to future work.
@@
where $C=0.75$ is a numerical factor and $\sigma_{\mathrm{eff}}$ is the effective cross section (Yang & Yu, 2022). For simplicity, we assume a constant cross section in this work. Figure 1 (right) shows the collapse time vs virial mass for the progenitors, where we have taken $\sigma/m=50~{}{\rm cm^{2}~{}g^{-1}}$. About one-third of the halos are expected to collapse within $10~{}{\rm Gyr}$. Since tidal stripping could speed up the onset of the collapse (Nishikawa et al., 2020; Sameie et al., 2020; Kahlhoefer et al., 2019; Yang & Yu, 2021; Zeng et al., 2022), we expect that more halos would be in the collapse phase after they evolve in the tidal field, and their overall mass would be reduced as well.
@@ 3.2 The Milky Way model
These parameters are motivated by the Milky Way mass model in McMillan (2016). Note that the stellar and disk density profiles in McMillan (2016) use exponential functions, which are challenging to implement in controlled N-body simulations due to the lack of analytical expressions for their corresponding potentials. Nevertheless, we have verified that the difference in the total potential remains within $2\%$ in the regions with $\sqrt{R^{2}+z^{2}}>10~{}{\rm kpc}$, which are most relevant for our simulated subhalo. Since the host halo is treated as a static potential, we neglect dark matter particle scatterings between the host halo and the subhalo. This approximation is well justified for velocity-dependent SIDM models with $\sigma/m\lesssim 1~{}{\rm cm^{2}~{}g^{-1}}$ at $V_{\rm max}\sim 200~{}{\rm km~{}s^{-1}}$ (Nadler et al., 2020a).
@@ 3.3 Orbital parameters
Figure 2: Left: evolution of the enclosed mass within $10~{}{\rm pc}$ for the CDM (blue), SIDM30 (amber), SIDM50 (orange), and SIDM100 (pink) subhalos. The horizontal dashed gray line denotes the inner mass within $10~{}{\rm pc}$ of the reference Hernquist profile, as shown in Figure 1 (middle). Middle: evolution of the bound mass for the simulated CDM and SIDM subhalos. Right: corresponding density profiles for the simulated CDM and SIDM subhalos at $t=10~{}{\rm Gyr}$, along with the initial NFW profile (black). The dotted blue line denotes a reconstructed density profile for the CDM subhalo using an analytical function proposed by Errani & Navarro (2021). The gray shaded region denotes the viable range for the GD-1 perturber, converted from Figure 6 of Bonaca et al. (2019), while the dashed gray line represents the reference Hernquist profile. The dashed cyan line represents the density profile of the globular cluster NGC 2419, modeled using the King profile from Baumgardt et al. (2009).
@@ 4 Results
Figure 2 (left) shows the evolution of the enclosed mass within the inner $r=10~{}{\rm pc}$ for CDM (blue), SIDM30 (amber), SIDM50 (orange), and SIDM100 (pink) subhalos. For CDM, the inner mass decreases monotonically due to tidal stripping. In contrast, for SIDM, the mass initially decreases sharply due to core expansion, followed by an increase as core collapse occurs. By $t\approx 10~{}{\rm Gyr}$, the inner mass of the SIDM subhalos is an order of magnitude higher than that of the CDM subhalo, aligning well with the reference Hernquist profile (horizontal line). Additionally, the collapse times are $t_{c}\sim 6~{}{\rm Gyr},~{}4~{}{\rm Gyr}$, and $2~{}{\rm Gyr}$ for SIDM30, SIDM50, and SIDM100, respectively, about a factor of two shorter than those estimated using Equation 2, which is calibrated for isolated halos. In a subhalo, tidal stripping reduces the velocity dispersion of dark matter particles from the intermediate to outer regions as a result of mass loss. Consequently, a negative “temperature” gradient—a necessary condition for the onset of core collapse—is more easily established compared to an isolated halo (Sameie et al., 2020).
@@
We note that the CDM subhalo has a small density core near the center. This is due to the resolution limit as $\ell=2.8\epsilon\approx 5.6~{}{\rm pc}$, although the simulated subhalo contains more than $6\times 10^{5}$ simulation particles at $t=10~{}{\rm Gyr}$. We fit the density profile using the analytical function $\rho(r)=\rho_{\rm cut}\exp(-r/r_{\rm cut})r_{\rm cut}/r$ from Errani & Navarro (2021), which is proposed to model a tidally stripped CDM halo. With $\rho_{\rm cut}\approx 1.2\times 10^{8}~{}M_{\odot}~{}{\rm kpc^{-3}}$ and $r_{\rm cut}\approx 0.22~{}{\rm kpc}$, we find a good fit for the region $r\gtrsim 10~{}{\rm pc}$, see Figure 2 (right, dotted blue). The fitted function has a cusp $\rho(r)\propto r^{-1}$ near the center and it provides a correction to the core due to the resolution limit. For the fitted profile, the enclosed mass within $10~{}{\rm pc}$ is $1.6\times 10^{4}~{}M_{\odot}$. However, even with this correction, the CDM subhalo remains insufficiently dense to explain the high density of the GD-1 perturber.
In Figure 2 (right), we also present the density profile of the globular cluster NGC 2419 (cyan), modeled using the King profile from Baumgardt et al. (2009). Interestingly, this profile closely resembles the density profile of the SIDM30 subhalo within $30~{}{\rm pc}$. This similarity is not coincidental, as the formation of globular clusters follows the same mechanism as the collapse of SIDM halos. This suggests that distinguishing between SIDM and globular cluster scenarios in explaining the GD-1 perturbation could be challenging. However, NGC 2419 itself cannot be the GD-1 perturber, as its orbit does not align with the perturbation (Bonaca et al., 2019). If the perturbation is caused by an undetected globular cluster that emits light, it could be identified in future astronomical surveys.
Furthermore, narrowing down the favored parameter space in the mass-size plane for the perturber would help us distinguish the two scenarios. For instance, if the perturber’s mass is further constrained to the range $10^{7}\textup{--}10^{8}~{}M_{\odot}$, the SIDM scenario would be favored, as globular clusters typically have masses below a few times $10^{6}~{}M_{\odot}$. Another intriguing possibility is that the perturber is an SIDM substructure hosting stars, as we will discuss later. It may have undergone significant tidal stripping, resulting in an ultrafaint dwarf with mass and structural properties similar to those of a massive globular cluster (Mau et al., 2020). Confirming this scenario would require detecting a stellar counterpart at the inferred location of the perturber. Distinguishing between these possibilities will require dedicated observational campaigns and detailed modeling efforts, making this an exciting avenue for future research.
@@ 5 Discussions and conclusion
The subhalo we used to demonstrate the SIDM scenario for the GD-1 perturber has an infall mass of $\approx 3\times 10^{8}~{}M_{\odot}$. Interestingly, this is near the upper limit on the peak mass of subhalos that host currently-observed satellite galaxies in the Milky Way (Jethwa et al., 2018; Nadler et al., 2020b). Thus, it remains an open question whether the perturber is a truly dark substructure, devoid of a galaxy. To further investigate detectability, we conducted additional simulations for the CDM and SIDM50 cases with live stellar particles, assuming a Plummer stellar profile with a scale radius of $0.3~{}{\rm kpc}$ and a total mass of $3.2\times 10^{4}~{}{M_{\odot}}$, motivated by hydrodynamical simulations of the Local Group (Fattahi et al., 2018). At $t=10~{}{\rm Gyr}$, the bound stellar masses are $1.8\times 10^{4}~{}M_{\odot}$ and $1.3\times 10^{4}~{}M_{\odot}$ for the CDM and SIDM50 cases, respectively, with the latter also exhibiting a steeper stellar density profile toward the central regions. These substructures fall into the category of ultrafaint dwarf galaxies and could potentially be detected in the near future through observations, e.g., with the Rubin Observatory (Ivezić et al., 2019; Drlica-Wagner et al., 2019).
More work is needed along these lines. For instance, the stellar-halo mass relation becomes increasingly steep in the ultrafaint regime and exhibits significant scatter (Fattahi et al., 2018), which must be taken into account. Additionally, since the Rubin Observatory can only detect objects in the southern hemisphere, it would be crucial to assess Rubin’s sky coverage in conjunction with the orbital information of the GD-1 perturber from Bonaca et al. (2020). We leave these investigations for future work. Furthermore, our scenario should also apply to smaller infall masses below $\sim 10^{8}~{}M_{\odot}$. Indeed, for SIDM models with large velocity-dependent cross sections, the population of core-collapsing (sub)halos increases as the mass decreases (e.g. Yang et al., 2023a; Nadler et al., 2023). Thus stellar streams like GD-1 can probe both population and density profile of core-collapsing subhalos even below the mass threshold for galaxy formation.